OpenStudy (anonymous):

Solve x2 + 8x - 48 = 0 by completing the square .

OpenStudy (karatechopper):

Show me what you got so far?

OpenStudy (anonymous):

what does completing the square mean ?

OpenStudy (anonymous):

You need to add a constant to both sides to create a "perfect square trinomial"

OpenStudy (anonymous):

sorry, what does that mean ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Like, x^2+2x+1 is a perfect square trinomial, as it can be condensed into (x+1)^2

OpenStudy (anonymous):

oh so like factor ?

OpenStudy (anonymous):

Sort of, you have to have the correct constant though or it won't work.

OpenStudy (anonymous):

(x-4) (x+12) =0

OpenStudy (anonymous):

like that ?

OpenStudy (anonymous):

That doesn't work. You have to complete the square by adding a constant to each side of the equals sign.

Directrix (directrix):

x2 + 8x - 48 = 0 x² + 8x = 48 x² + 8x + _?__ = 48 + __?__ The object is to force the left side of the equation to be a perfect square trinomial.

Directrix (directrix):

x² + 8x + _?__ = 48 + __?__ Take half of the coefficient of the linear x term (it is 8) --> 4 Square it. 4² = 16 x² + 8x + _16__ = 48 + __16__

Directrix (directrix):

@koreanhugg How does x² + 8x +16 factor?

Directrix (directrix):

@koreanhugg is "is offline"

OpenStudy (anonymous):

omg sorry ! my internet was being dumb . .

Directrix (directrix):

Let's pick up where you left off: @koreanhugg How does x² + 8x +16 factor?